A transparent photographic slide is placed in fron
Answer #1

a)

if image is real then image distance is +ve, formed on the opposite side of the lens

(1/f) = (1/dₒ) + (1/dᵢ)

(1/2.41) = (1/dₒ) + (1/13.2)

1/dₒ = 0.33918

dₒ = 2.95 cm [object is between F and 2F]

b)

image is virtual:

(1/2.41) = (1/dₒ) + (1/-13.2)

(1/dₒ) = 0.490

dₒ = 2.038 cm [object is between lens and F(focal point)]